OpenStudy (anonymous):

lim x->0 of (1-ax)^(1/x)

terenzreignz (terenzreignz):

Is this e?

terenzreignz (terenzreignz):

lol, too good to be true

OpenStudy (anonymous):

what do you mean by is this e?

terenzreignz (terenzreignz):

I meant the constant e, but I don't want to conclude that, I have a bad feeling about it ;)

OpenStudy (anonymous):

the question does not have e in it but i believe the answer will, i think i have to natural log it or something. i really am not sure.

OpenStudy (anonymous):

answer is 1

OpenStudy (anonymous):

msamido, can you please show the steps?

terenzreignz (terenzreignz):

Uhh, this limit seems like one indeterminate form

OpenStudy (anonymous):

yeah. i have to turn it into a fraction and apply the la'hopital rule. but , i dont know how.

OpenStudy (anonymous):

\[\huge \lim_{x \rightarrow 0} (1-ax)^{1/x}\] Indeterminate form of type 1^infinity. Transform using: \[\huge \lim_{x \rightarrow 0} (1-a x)^{1/x} = e^{\lim_{x \rightarrow 0} \frac{ln(1-a x)}{x}}\] into: \[\huge e^{\lim_{x \rightarrow 0} \frac{ln(1-a x)}{x}} \] Indeterminate form of type 0/0, so use L'Hospital's rule \[\huge e^{\lim_{x \rightarrow 0} \frac{a}{a x-1}} \] Factor out constants: \[\huge e^{a (\lim_{x->0} \frac{1}{a x-1})} \] The limit of a quotient is the quotient of the limits: \[\huge e^{\frac{a}{lim_{x->0} (a x-1)}} \] The limit of a x-1 as x approaches 0 is -1: \[\huge \frac{1}{e^a}\ = e^{-a}]

OpenStudy (anonymous):

Ugh syntax... The limit of a x-1 as x approaches 0 is -1: \[\huge \frac{1}{e^a}\ = e^{-a}\]

OpenStudy (anonymous):

This was kind of a pain to type up with all those {}'s :-P Are you lost on any step? I don't think I have any more typos, but I've been trying to weed those out for the past few minutes, so hey there could be more

OpenStudy (anonymous):

lim x->0 (1-ax)^(1/x) introducing natural logarithm.. =(1/x)ln(1-ax) =lim x->o((1/x)(ln(1-ax)) when substituting u will get indeterminant of the form 0/0 them we apply L'HOSPITALS RULE =limx->0(1/(1-ax)) ax will approach 0 when x -> 0 =1/1 =1

OpenStudy (anonymous):

YES. I GOT ITTTTTT! :D thank you sooooooooooo much!

OpenStudy (anonymous):

thanks!!